OpenStudy (anonymous):

An ideal gas is in equilibrium at initial state with temperature T=137oC, pressure P = 0.75Pa and volume V = 0.75 m3. If there is a change in state in which the gas undergoes an isothermal process to a final state of equilibrium during which the volume is doubled. Calculate the temperature and pressure of the gas at this final state.

OpenStudy (anonymous):

WKT P*V=n*R*T Since it is an isotheremal process therefore, n, R , T are constant Therefore temperature remains 137oC Use (P_i)*(V_i) =(P_f)*(V_f) (0.75)(0.75)=(P_f)*(1.50) P_f = 0.35 Pa

OpenStudy (anonymous):

P=pressure V=volume n=no of moles R=gas constant T=temperature

OpenStudy (anonymous):

@shivam_bhalla what about the final temperature

OpenStudy (anonymous):

Sorry P_f = 0.365 Pa

OpenStudy (anonymous):

Since it is an isothermal process, therefore temperature remains constant